# AgriFM: A Multi-source Temporal Remote Sensing Foundation Model for Crop Mapping

### v2

@@ 6 Winter Wheat Mapping in Asia
Several key observations emerge from the comparative analysis. First, Swin-based methods consistently dominate the rankings, with PIS (74.47% F1) and Swin-scratch (74.23% F1) already surpassing all CNN and ViT variants. This 6-8% F1 advantage over ViT models particularly highlights the importance of hierarchical feature extraction when working with low-resolution inputs. Notably, the standard CNN achieves better performance (70.17% F1) than more complex 3DCNN (67.69% F1) and CNN-LSTM (68.27% F1) architectures, suggesting that simple spatial feature extractors may be more effective than sophisticated spatiotemporal designs for this specific task.
The superior performance of AgriFM can be attributed to two synergistic factors. First, its MODIS-specific pretraining captures unique phenological patterns critical for winter wheat identification - an advantage missing in other models. Second, the architecture’s multi-source design enables exceptional feature extraction from low-resolution data, evidenced by its best-in-class recall (78.74%), indicating superior detection of true wheat areas. This combination proves especially valuable for winter wheat mapping, where the crop’s distinct growth stages create temporally identifiable signatures in MODIS data.
In addition, while winter wheat’s strong phenological characteristics make it relatively distinguishable from other crops, the performance gaps between methods are larger than those observed in rice mapping (Table 6). This accentuates our model’s specialized capabilities - the 8.8% F1 improvement over ViT models and 5.2% over CNNs demonstrate AgriFM’s particular aptitude for leveraging subtle temporal cues in coarse-resolution data.
Table 7: Comparison results on winter wheat mapping in Asia with different methods.
@@
AgriFM 72.34 78.74 75.40 85.53 88.47 86.93 97.11
Figure 12: Visual comparison of winter wheat mapping in Asia.
@@ 7 Discussions
To further validate the robustness and versatility of our proposed framework, we conduct more systematic analyses addressing critical operational aspects of crop mapping tasks: temporal adaptability, and data efficiency. These experiments are designed as follows.
  • Quantify the model’s capacity to handle variable-length temporal sequences—a fundamental requirement for crop mapping across diverse phenological cycles and cloud cover conditions. We systematically evaluate its performance on early-season crop mapping using temporally downsampled input sequences. Specifically, we down-sample the original Sentinel-2 time series into fixed-interval sequences of 8-, 12-, and 16-step configurations. This simulates scenarios with varying observational frequencies while preserving phenological relevance.
  • Evaluate practical applicability under constrained label availability scenarios (10%, 20%, 33%, 50% subsets). Critically, we compare against non-pretrained variants of our framework (random initialization) to quantify the benefits of foundation model pretraining. This demonstrates the value of large-scale self-supervised learning in label-scarce agricultural settings.
Table 8: Evaluation of the model’s capacity to handle variable-length temporal sequences (8-, 12-, 16-, and 24-step sequences).
Sequence length Precision Recall F1 OA
8-step sequences 69.05 45.63 50.50 74.55
12-step sequences 70.07 48.72 54.87 76.14
16-step sequences 70.50 50.07 56.45 76.23
24-step sequences 68.63 53.88 59.88 77.43
Table 9: Evaluate practical applicability under constrained label availability scenarios (10%, 20%, 33%, 50% subsets).
Training subsets Model Precision Recall F1 OA
10% wo/ pre-training 53.21 33.87 38.28 71.45
w/ pre-training 59.31 32.56 37.65 70.67
20% wo/ pre-training 57.11 41.36 46.77 74.28
w/ pre-training 64.36 42.07 48.11 73.92
33% wo/ pre-training 65.54 45.04 50.43 75.31
w/ pre-training 62.10 48.09 52.85 75.57
50% wo/ pre-training 65.40 47.75 54.06 75.70
w/ pre-training 65.42 52.17 57.63 76.26
100% wo/ pre-training 70.43 48.35 54.82 76.20
w/ pre-training 68.63 53.88 59.88 77.43
Figure 13: Systematic analyses addressing critical operational aspects of crop mapping tasks: temporal adaptability and data efficiency.
As evidenced by Table 8, the framework demonstrates progressive performance gains with extended temporal sequences, attaining an F1 score of 59.88% and overall accuracy (OA) of 77.43% for 24-step sequences. A critical observation is the pronounced improvement in recall, which rises from 45.63% (8-step) to 53.88% (24-step)—a 17.5% relative increase—indicating enhanced discrimination of crop-specific phenological patterns as temporal contexts expand. This aligns with our line chart visualizations from Figure 13 a), where recall exhibits a near-linear correlation with sequence length, suggesting that longer time series better resolve phenological ambiguities. Notably, even under sparse 8-step inputs, the model retains 84.3% (50.50 vs. 59.88) of its peak F1 performance, underscoring its effectiveness in leveraging both spectral-spatial patterns and partial phenological cues. Interestingly, while precision peaks at 70.50% for 16-step sequences before slightly declining, OA consistently improves (+3.88% from 8- to 24-step). These results validate the framework’s dual capacity: preserving spatial fidelity under temporal sparsity while capitalizing on extended sequences to capture growth-stage dynamics.
The label efficiency analysis (Table 9 and Figure 13 b)) underscores the practical value of our foundation model pretraining strategy. As the fine-tuning dataset scales from 10% to 100%, the pretrained variant demonstrates consistent performance gains, achieving a 57.63% F1 score with only 50% fine-tuning data—surpassing the non-pretrained model’s full-data performance (54.82%). This highlights the pretraining framework’s ability to extract transferable spatiotemporal representations, enabling effective adaptation to downstream tasks with limited annotations. Notably, while the non-pretrained model plateaus beyond 50% fine-tuning data ($\Delta$F1=+0.76from 50% to 100%), the pretrained variant continues to improve ($\Delta$F1=+2.25%), suggesting superior utilization of additional labeled samples.
At the 10% fine-tuning level, the pre-trained model exhibits a precision-recall imbalance (59.31% precision vs. 32.56% recall), with recall nearly matching the non-pretrained counterpart (32.56% vs. 33.87%). This phenomenon likely stems from insufficient fine-tuning data to fully adapt the pretrained features for distinguishing subtle inter-class variations, causing the model to prioritize precision over recall. However, as fine-tuning data increases to 20%, the pretrained model establishes a decisive advantage (+1.34% F1), demonstrating its capacity to leverage pretrained knowledge for improved generalization. Notably, the pretrained model achieves comparable performance to its non-pretrained counterpart using significantly less data: at just 33% fine-tuning data (F1=52.85%), it matches the non-pretrained model’s full-data performance (F1=54.82%). This demonstrates the strong transferability of pretrained representations, enabling effective deployment in regions with limited labeled data.

### v3

@@ 6 Winter Wheat Mapping in Asia
Several key observations emerge from the comparative analysis. First, Swin-based methods consistently dominate the rankings, with PIS (74.47% F1), GFM (74.33% F1) and Swin-scratch (74.23% F1) already surpassing all CNN and ViT variants. This 4-8% F1 advantage over ViT models particularly highlights the importance of hierarchical feature extraction when working with low-resolution inputs. Notably, the standard CNN achieves better performance (70.17% F1) than more complex 3DCNN (67.69% F1) and CNN-LSTM (68.27% F1) architectures, suggesting that simple spatial feature extractors may be more effective than sophisticated spatiotemporal designs for this specific task.
The superior performance of AgriFM can be attributed to two synergistic factors. First, its MODIS-specific pretraining captures unique phenological patterns critical for winter wheat identification - an advantage missing in other models. Second, the architecture’s multi-source design enables exceptional feature extraction from low-resolution data, indicating superior detection of true wheat areas. This combination proves especially valuable for winter wheat mapping, where the crop’s distinct growth stages create temporally identifiable signatures in MODIS data.
In addition, while winter wheat’s strong phenological characteristics make it relatively distinguishable from other crops, the performance gaps between methods are larger than those observed in rice mapping (Table 6). This accentuates our model’s specialized capabilities - the 4.4% F1 improvement over ViT models and 5.68% over CNNs demonstrate AgriFM’s particular aptitude for leveraging subtle temporal cues in coarse-resolution data.
Table 7: Performance comparison on the winter wheat mapping task (%). The positive column means metrics of winter wheat identification, while the average columns means the average metric of both background and winter wheat.
@@
AgriFM 75.56 76.14 75.85 87.07 87.34 87.20 97.27
Figure 12: Visual comparison of winter wheat mapping results in Asia from test dataset in 2021. Only the best-performing methods from each task are shown to ensure clarity in comparison. Winter wheat pixels are shown in green. Representative regions with notable differences are highlighted with red bounding boxes for detailed comparison.
@@ 7 Discussions
7.1 Data-Efficiency Evaluation of Foundation Models with Varying Fine-tuning Samples
To comprehensively evaluate the data efficiency of different foundation models, we conducted extensive experiments across five agricultural mapping tasks with progressively reduced training data ratios (from 100% to 5%). This analysis aims to assess the robustness and generalization capability of foundation models when facing limited annotated samples, which is a common challenge in real-world agricultural applications. The experimental setup follows the standard fine-tuning protocol where models pre-trained on large-scale datasets are adapted to downstream tasks with varying amounts of supervised data (Marsocci et al., 2025). Table 8 presents the detailed performance comparison across all training ratios and tasks, while Figure 13 illustrates the performance trends through line charts, with the first five subplots corresponding to individual tasks and the final subplot showing the averaged performance across all tasks.
The experimental results demonstrate that our proposed AgriFM consistently achieves superior performance across nearly all tasks and training ratios, confirming its exceptional capability in handling data-scarce scenarios. For agricultural land mapping and field boundary delineation tasks, AgriFM maintains a substantial performance advantage throughout all data regimes, with improvements of 2-5% over the second-best methods. This consistent superiority suggests that the representations learned by AgriFM are particularly well-suited for these fundamental agricultural mapping tasks, even when fine-tuning samples are severely limited.
However, a distinct pattern emerges for the more complex agricultural land use/cover mapping task, where all models experience rapid performance degradation as training data decreases. While AgriFM still achieves the best performance in this challenging scenario, the dramatic performance drop across all models underscores that complex fine-grained classification tasks inherently require substantial annotated data to maintain satisfactory performance levels. This observation highlights an important limitation of current foundation models and suggests that for such detailed classification problems, having adequate training samples remains crucial.
Table 8: Performance comparison of foundation models across varying training data ratios (Metrics: mF1 for Agri. land cover / land use mapping, F1 on positive class for other tasks). Best performance for each ratio is bolded.
Model 100% 50.0% 33.0% 25.0% 20.0% 10.0% 5.0%
Agricultural land mapping
Prithvi 73.69 69.51 69.23 69.30 66.38 66.07 63.61
Gelileo 75.73 75.62 73.41 68.08 69.24 68.12 65.41
SatMAE 76.80 75.22 74.70 72.66 70.49 69.74 67.69
SMARTIES 74.76 69.26 69.60 69.63 66.91 64.63 61.16
PIS 79.01 78.00 75.95 75.52 75.00 74.15 72.33
GFM 81.06 77.24 75.25 75.68 75.35 74.15 72.06
AgriFM 83.09 81.66 80.96 80.57 80.31 76.73 74.27
Field boundary delineation
Prithvi 53.45 55.54 49.52 56.62 56.04 54.27 49.21
Gelileo 62.51 56.01 56.47 55.91 55.30 55.26 52.18
SatMAE 62.50 59.89 57.91 55.83 54.95 53.96 52.58
SMARTIES 59.92 56.86 55.98 54.95 53.90 53.02 52.89
PIS 71.20 69.81 66.57 65.03 62.22 63.17 59.70
GFM 72.38 68.91 66.60 65.30 64.40 62.94 62.60
AgriFM 76.27 73.55 71.86 70.83 69.91 68.38 66.36
Agricultural land use / land cover mapping
Prithvi 42.10 37.91 35.01 35.40 34.91 30.26 25.28
Gelileo 45.12 40.65 40.27 37.20 35.87 32.09 26.40
SatMAE 46.10 42.13 41.10 37.83 35.59 31.61 25.22
SMARTIES 47.65 40.48 39.13 38.30 35.64 32.02 25.42
PIS 54.51 46.02 46.49 43.65 41.44 35.72 29.22
GFM 57.75 49.29 47.90 44.60 42.52 35.01 30.21
AgriFM 60.49 52.00 49.51 45.67 42.64 35.47 30.33
Paddy rice mapping
Prithvi 82.64 81.61 80.74 80.10 79.99 78.92 77.01
Gelileo 84.79 82.27 81.44 80.90 80.75 79.17 76.87
SatMAE 83.22 82.28 81.30 80.46 80.70 79.00 76.68
SMARTIES 84.10 82.03 77.31 80.12 79.94 78.99 75.29
PIS 86.49 85.19 84.01 83.36 83.08 81.75 78.63
GFM 86.31 84.76 83.61 83.36 83.09 80.92 76.90
AgriFM 86.97 86.06 85.31 84.76 84.63 82.70 79.53
Winter wheat mapping
Prithvi 66.61 57.88 62.73 58.31 49.83 54.18 49.92
Gelileo 71.45 63.23 65.82 60.42 60.81 56.40 50.55
SatMAE 67.17 64.27 65.14 60.49 61.74 57.14 50.15
SMARTIES 65.34 57.35 58.94 57.44 56.02 45.32 36.48
PIS 74.47 73.34 73.93 71.55 71.49 69.17 65.80
GFM 74.33 73.19 73.82 70.96 71.73 69.81 66.10
AgriFM 75.85 75.17 74.70 73.03 73.81 72.16 68.82
Figure 13: Performance trends of foundation models across different training data ratios. The first five subplots corresponding to individual tasks and the final subplot showing the averaged performance across all tasks.
For specific crop mapping tasks including paddy rice and winter wheat, AgriFM not only achieves the best performance under full data conditions but also demonstrates particularly significant advantages in low-data regimes. The performance gap between AgriFM and other foundation models widens notably when training data is reduced to 20% or lower, indicating that our model’s pre-training strategy effectively captures crop-specific phenological patterns that transfer well even with minimal fine-tuning samples. This characteristic is especially valuable for practical applications where collecting large annotated datasets for specific crops may be challenging or costly.
In summary, these comprehensive experiments validate AgriFM’s superior data efficiency and robustness across diverse agricultural mapping tasks. The consistent performance advantage, particularly in low-data scenarios, demonstrates that our method learns more transferable representations during pre-training that require less supervised data for effective adaptation to downstream tasks. These findings position AgriFM as a practical solution for agricultural monitoring applications where annotated data may be limited, while also highlighting the ongoing challenges in complex fine-grained classification scenarios.
7.2 Ablation Study on Key Components of AgriFM
To systematically evaluate the contribution of each component in our proposed AgriFM framework, we conduct comprehensive ablation studies focusing on both pre-training strategies and architectural designs. The experiments are designed to isolate and quantify the impact of four key pre-training elements: (1) the mean-teacher framework for handling label noise in land cover products, (2) land cover fraction supervision during pre-training, (3) multi-source data integration across MODIS, Landsat, and Sentinel-2, and (4) multi-temporal sampling with variable sequence lengths (3-32 frames). Additionally, we analyze the effectiveness of our synchronized temporal downsampling strategy in balancing computational efficiency and performance. For the multi-source ablation, when disabled, we pre-train three separate models using individual data sources and fine-tune each downstream task with the model corresponding to its input data source. Similarly, for the multi-temporal ablation, the disabled setting employs the conventional fixed 16-frame input commonly used in video foundation models.
The results presented in Table 9 demonstrate the progressive improvement achieved by incorporating each component into our framework. Starting from a baseline without mean-teacher and land cover supervision (first row), we observe that adding land cover supervision provides mixed benefits, showing improvements in some tasks while slightly decreasing performance in others. This suggests that while land cover supervision provides valuable semantic guidance, its effectiveness can be limited by label noise in the absence of proper regularization. The incorporation of multi-source data consistently enhances performance across most tasks, particularly benefiting agricultural land mapping and boundary delineation. Notably, the mean-teacher framework emerges as a crucial component, substantially boosting performance across all tasks when combined with other elements. The complete AgriFM configuration with all components achieves the best overall performance, with particularly remarkable improvements in winter wheat mapping (76.65% F1) and boundary delineation (76.27% F1), demonstrating the synergistic effect of integrating these complementary design choices.
Table 9: Ablation study on pre-training strategies. MT: Mean-Teacher structure; LC: Land Cover fraction supervision; Multi-source: Using multiple satellite sources during pre-training; Multi-temporal: Variable sequence length input (3-32 frames) during pre-training. Best performance for each task is bolded.
MT LC Multi-source Multi-temporal
Agri. land
mapping
Boundary
delineation
Agri. land
use/cover
Paddy
rice
Winter
wheat
$\times$ $\times$ 81.96 73.70 59.73 86.53 73.69
$\times$ 81.42 74.35 59.82 86.05 74.18
$\times$ $\times$ 82.70 75.22 59.93 86.09 74.08
$\times$ 81.90 73.78 59.88 86.04 75.40
83.09 76.27 60.49 86.97 76.65
Table 10 presents a focused analysis of our synchronized temporal downsampling strategy, comparing both performance and computational requirements against a baseline without temporal downsampling. The metrics include: time (total training hours), Mem/b (memory consumption per batch in GB, with batch size indicated after the slash), and FLOPs (floating point operations in GigaFLOPs). While the non-downsampling variant achieves marginally higher performance in agricultural land mapping (83.53% vs 83.09%), it suffers significant performance degradation in agricultural land use / land cover mapping (57.58% vs 60.49%). More importantly, the computational advantages of our approach are substantial, reducing training time by approximately 40-45%, memory consumption by 15-33%, and FLOPs by 60-70% across different tasks. This efficiency gain is particularly valuable for large-scale agricultural monitoring applications where computational resources are often constrained.
These ablation studies collectively demonstrate that each component in AgriFM contributes meaningfully to its overall effectiveness, with the full configuration achieving an optimal balance between performance and efficiency. The synchronized temporal downsampling strategy emerges as particularly crucial, enabling efficient processing of long temporal sequences while maintaining competitive performance across diverse agricultural mapping tasks.
Table 10: Performance and computational efficiency of synchronized temporal downsampling. Metrics: time (training hours), Mem/b (memory consumption in GB per batch with batch size), FLOPs (Giga floating-point operations). Best performance for each task is bolded.
Agri. land mapping Boundary delineation Agri. land use / land cover
Method F1 Time
Mem/b
(GB/bs)
FLOPs F1 Time
Mem/b
(GB/bs)
FLOPs mF1 Time
Mem/b
(GB/bs)
FLOPs
w/o downsampling 83.53 17h 36.2G/2 804G 76.16 17h 36.2G/2 580G 57.58 16h 36.3G/3 580G
AgriFM 83.09 10h 30.6G/4 256G 76.27 10h 30.6G/4 256G 60.49 9.5h 24.5G/4 220G
7.3 Cross-Temporal and Cross-Source Generalization Analysis
To address the critical aspect of foundation model generalization capabilities, we conducted comprehensive experiments to validate AgriFM’s generalization capabilities across temporal dimensions and sources with different spatial resolutions. These experiments assess the model’s adaptability to varying input configurations, including different temporal sequence lengths and diverse satellite data sources with varying spatial resolutions. The cross-temporal analysis evaluates performance stability across different temporal contexts, while the cross-spatial analysis examines the model’s ability to leverage multi-resolution satellite data for enhanced agriculture mapping.
The temporal generalization analysis examines model performance with input sequence lengths ranging from 4 to 24 frames. As demonstrated in Table 11 and Figure 14, AgriFM consistently outperforms all competing foundation models across all temporal configurations. Notably, while most models exhibit performance degradation with shorter sequences, AgriFM maintains robust performance even with 12-frame inputs (59.00% mF1), demonstrating only a marginal 1.49% decrease from the 24-frame configuration. This temporal robustness stems from our synchronized spatio-temporal downsampling strategy, which enables effective feature extraction regardless of input sequence length.
Table 11: Performance comparison (mF1 scores) on agricultural land use/cover mapping with varying input sequence lengths (4, 8, 12, 16, 24 frames). Best performance for each sequence length is bolded.
Model 24 frames 16 frames 12 frames 8 frames 4 frames
Prithvi 42.10 43.01 41.60 38.93 39.85
Gelileo 45.12 43.52 42.60 41.81 39.75
SatMAE 46.10 45.34 44.70 43.49 39.69
SMARTIES 47.65 44.15 43.20 41.52 39.50
PIS 54.51 53.32 49.40 47.13 44.11
GFM 57.75 53.23 53.00 48.46 46.24
AgriFM 60.49 59.87 59.00 53.62 48.08
Figure 14: Performance comparison (mF1 scores) on agricultural land use/cover mapping with varying input sequence lengths (4, 8, 12, 16, 24 frames).
The cross-source generalization analysis, presented in Table 12, reveals insightful patterns about AgriFM’s ability to leverage diverse satellite data sources. Sentinel-2 data (10m resolution) serves as our high-resolution baseline, providing the highest individual performance across most tasks due to its superior spatial and temporal resolution. However, its temporal coverage is limited to post-2017 data. Landsat 8/9 data (30m resolution) with 12 monthly frames shows consistently lower performance than Sentinel-2, with particularly pronounced degradation in complex tasks like agricultural land use/cover mapping (33.59% vs 60.49% mF1).
The integration of MODIS data with its comprehensive 44-frame annual coverage demonstrates complementary benefits. For cropland mapping, the Landsat+MODIS combination (83.14% F1) surpasses Sentinel-2 alone (83.09% F1), with similar improvements observed across other tasks. This suggests that MODIS’s dense temporal sampling provides valuable phenological information that compensates for its coarser spatial resolution. Furthermore, the fusion of Sentinel-2 and Landsat data yields substantial improvements in agricultural land mapping (85.61% vs 83.09% F1) and field boundary delineation (78.39% vs 76.27% F1), indicating that the complementary characteristics of these sensors enhance spatial feature extraction.
However, for the more complex agricultural land use/cover mapping task, the benefits of data fusion are more nuanced. While Sentinel-2 alone achieves 60.49% mF1, the addition of Landsat data provides only marginal improvement (60.57% mF1), suggesting that Landsat’s limited additional information has minimal impact on fine-grained classification performance. This pattern highlights that while multi-source integration generally benefits agricultural mapping tasks, its effectiveness is task-dependent, with complex classification problems requiring the high-quality features provided by premium data sources like Sentinel-2.
These comprehensive experiments demonstrate that AgriFM possesses strong cross-temporal and cross-source generalization capabilities, effectively leveraging diverse temporal sequences and multi-resolution satellite data to enhance agricultural mapping performance across various tasks and configurations.
Table 12: Performance across different satellite data sources with varying spatial resolutions (Sentinel-2: 10m&20m; Landsat 8/9: 30m; MODIS: 250m&500m) and temporal configurations. Numbers in data source columns indicate temporal frame counts.
Task
Sentinel-2
frames
Landsat
frames
MODIS
frames
Precision Recall F1
Agri. land mapping 32 84.31 81.90 83.09
12 83.54 81.78 82.65
32 12 86.18 85.05 85.61
12 44 71.74 84.58 83.14
Field boundary delineation 32 75.11 77.47 76.27
12 72.78 73.45 73.11
32 12 77.91 78.88 78.39
12 44 73.51 74.14 73.82
Agri. land use/cover mapping 24 68.34 54.97 60.49
12 49.22 27.86 33.59
24 12 68.37 55.08 60.57
12 44 55.26 28.55 34.95
Paddy rice mapping 5 84.03 90.12 86.97
5 44 83.61 91.63 87.44
7.4 Limitation and future directions
While the proposed AgriFM demonstrates strong performance across multiple agricultural mapping tasks, we acknowledge several limitations that point toward valuable future research directions. This part focuses on two primary aspects: the data alignment strategy employed during pre-training and the generalizability beyond the current application scope.
Data alignment strategy: Our method employs a non-paired multi-source learning strategy, which differs from conventional approaches that require spatial alignment across different satellite sensors. This design choice is supported by the use of land cover fraction supervision, which serves as a semantic bridge that enables feature-space alignment across varying spatial resolutions and sensor characteristics. The land cover fractions provide a consistent learning target that transcends the specific geometric properties of individual sensors, allowing the model to learn robust representations that are invariant to the particularities of each data source.
While our non-paired approach has demonstrated effectiveness in learning transferable representations, future work could explore the complementary benefits of paired multi-source data. A carefully constructed fully-aligned dataset with land cover supervision would enable more direct comparisons between alignment strategies and potentially reveal additional insights about cross-sensor relationships. Such investigations could further advance our understanding of optimal strategies for integrating heterogeneous remote sensing data in foundation model development.
Versatility of decoder: The current architecture of decoder, though unified across different mapping tasks, still requires task-specific fine-tuning and may not generalize effectively to fundamentally different problem paradigms beyond dense spatial prediction. Specifically, our decoder is not optimized for tasks such as site-based regression, point prediction, or other non-spatial learning objectives that diverge from the pixel-wise mapping framework. This limitation stems from our design focus on addressing the specific challenges of agricultural land monitoring, where the output consistently takes the form of spatial maps derived from satellite image sequences.
Looking forward, several promising directions emerge for enhancing decoder generalizability. Future work could explore the development of truly task-agnostic decoders through multi-task learning frameworks that simultaneously handle diverse agricultural monitoring objectives. Additionally, investigating dynamic decoder architectures that can adapt their structure based on task requirements presents an exciting research avenue. The integration of prompt-based mechanisms or task-conditioning approaches could further improve the decoder’s flexibility across different agricultural applications. Another valuable direction involves extending the decoder’s capabilities beyond traditional classification tasks to encompass regression-based agricultural monitoring, such as yield prediction or biophysical parameter estimation, while maintaining the architectural efficiency that characterizes our current design.
Spatial generalization of remote sensing foundation models: The geographical scope of our pre-training data remains limited compared to the global coverage of modern satellite systems. The fundamental question of how to optimize pre-training strategies for maximum geographical transferability needs deeper investigation. Unlike other domains where data can be considered independent samples, remote sensing observations are intrinsically spatial-dependent and constrained to Earth’s surface. This unique characteristic raises important questions about whether including specific geographical regions during pre-training systematically improves downstream performance in those same region.
The development of truly global foundation models presents both unprecedented opportunities and substantial challenges. The construction of a comprehensive global temporal dataset would require huge storage and computing capabilities. However, such investment could yield significant returns, potentially enabling robust zero-shot generalization through integration with multi-modal data sources like existing global products and language models. The parallel with scaling laws in language models suggests that increased data coverage and model capacity may systematically improve performance, though the unique spatial constraints of geographical data may introduce fundamentally different scaling behaviors that merit dedicated investigation.